OpenStudy (anonymous):

integral cos(3x)cos(4x) dx

OpenStudy (anonymous):

i dont even no where to begin

OpenStudy (anonymous):

the only way i think of to handle this is to use Euler's Formula's.\[\cos(3x)=\frac{e^{3ix}+e^{-3ix}}{2}\] \[\cos(4x)=\frac{e^{4ix}+e^{-4ix}}{2}\] Multiplying and simplifying should make the problem easier to handle.

OpenStudy (anonymous):

i havnet even come across that yet

OpenStudy (anonymous):

then maybe using sum and difference formulas? cos of two angles and whatnot? it would be really easy with Eulers though =/

OpenStudy (anonymous):

can you show me the first step with the sum and diffrence formulas

myininaya (myininaya):

equation1 : \[\cos(u+v)=\cos(u)\cos(v)-\sin(u)\sin(v)\] equation 2 : \[\cos(u-v)=\cos(u)\cos(v)+\sin(u)\sin(v)\] adding equation1 to equation2 we get \[\cos(u+v)+\cos(u-v)=2\cos(u)\cos(v)\] now multiply both sides by 1/2 \[\cos(u)\cos(v)=\frac{1}{2}(\cos(u+v)+\cos(u-v))\]

myininaya (myininaya):

\[\int\limits_{}^{}\cos(3x)\cos(4x)dx=\frac{1}{2}\int\limits_{}^{}(\cos(3x+4x)+\cos(3x-4x))dx\] \[=\frac{1}{2}\int\limits_{}^{}(\cos(7x)+\cos(-x)) dx=\frac{1}{2}\int\limits_{}^{}(\cos(7x)+\cos(x)) dx\]

myininaya (myininaya):

\[=\frac{1}{2}(\frac{1}{7}\sin(7x)+\sin(x))+C\]